OpenStudy (anonymous):

Factor completely. 4x^2+14x+12

OpenStudy (anonymous):

= 2(2x^2 + 7x +6) = 2(2x+3)(x+2)

OpenStudy (anonymous):

2(x+2)(2x+3)

OpenStudy (anonymous):

thanks :)